# What is a Digital Twin Anyway? Deriving the Definition for the Built Environment from over 15,000 Scientific Publications

### v1

@@ 3.1 Sample Development
This study is built on two main data sources. The first consists of 1) full-text articles, while the second involves a 2) expert survey which is used to validate the results from a Built Environment expert perspective. From these two datasets, we have derived: the 3) DT definitions dataset and the 4) DT components dataset. In this section, we overview our methodology of collecting the sample data and extracting the definitions’ and components’ datasets that are used for analysis.
@@ 3.1.1 Full-text articles dataset
Full-text samples were acquired using an Application Programming Interface (API) provided by Elsevier [^5] https://dev.elsevier.com/, which provides interfaces to Scopus, Science Direct, and other Elsevier products. We chose to focus on the Elsevier API because it is one of the largest publishers of scientific literature and provides a comprehensive database of peer-reviewed literature across myriads of fields, thus well capturing the body of knowledge without bias. We used the API to search for articles containing the following terms: “Digital Twin”, “City Digital Twin”, “Urban Digital Twin”, “Smart City”, and “Building Digital Twin” in the title, abstract, or keywords. The search was conducted in English and covered articles published between 2000 and 2024. The search returned a total of 15,353(Figure 6). The dataset was also collected from various sources (Figure 5), notably the top are from manufacturing (e.g, IFAC Journal of Systems and Control, Journal of Manufacturing Systems, Procedia Manufacturing); and the Built Environment (e.g, Automation in Construction, Transportation Research Procedia, Journal of Building Engineering, Energy and Buildings, and Building and Environment). The full texts of which were then downloaded and stored in a database for further analysis alongside their metadata including the publication type (full-text article, abstract, chapter, editorial, etc.), subjects, authors, date of publication, and keywords.
@@
Figure 6: Distribution of subjects of the selected dataset.
@@ 3.1.2 The expert survey dataset
The expert survey invited a structured panel of experts, who have domain expertise in digital twins related to the built environment, aiming to generate a consensus on the definition of digital twins. A total of 52 international experts from different sectors, such as government, industry (e.g. companies, consultancies), and non-government organizations (NGOs), were selected to build the panel, contributing with diverse backgrounds and experiences in this topic. Through an online questionnaire, all participants were asked to identify a digital twin based on their understanding and practice.
@@ 3.1.3 The definitions dataset
This pattern extracts all sentences that match “digital twin/s (is/are/can be/could be) (defined as/described as/ characterized by) …”. This method has proven to be efficient in scanning and extracting a large number of mentions and descriptions of DTs. However, many of the extracted matches do not qualify as definitions. For example, sentences such as “Digital Twin is a revolutionizing technology in the industry 4.0 era” and “Digital Twin is defined as a new paradigm in simulation” are merely characterizations, contextualizations, or portrayals of DTs, rather than definitions.Therefore, we applied Natural Language Processing (NLP) and Large Language Models (LLMs) to filter the dataset, aiming to address two main issues: Incomplete and Fuzzy Duplicated definitions.
@@
To perform the filter, we first excluded all the incomplete definitions using LLMs, namely, the Generative Pre-trained Transformer (GPT4) model developed and hosted by OpenAI. Then, we manually validated the results.
Secondly, to overcome the FDD issue, we first clustered the definitions using sentence embeddings Le and Mikolov (2014) and k-Means clustering into three iterations of clustering 400, 100, and 50 clusters. The Sentence Embeddings first convert all the definitions into a lower dimensional vector where the position of each definition in the vector space represents its semantic meaning. By doing this, the closer the definitions in the vector space, the more semantically similar these sentences are to each other. Each sentence is represented as a feature vector $\vec{X}_{d}=[x_{0},x_{1},x_{2},\ldots,x_{n}]$ where $X_{d}$ is the definition sentence vector. We used the cosine similarity distance to measure the similarity between different definitions in the definition dataset. The cosine similarity distance between two vectors $\mathbf{X}$ and $\mathbf{Y}$ is given by:
@@
After clustering the definitions into groups, we applied the Fuzzy Matching algorithm within each group to figure out how each pair of definitions matches. By doing so, we semi-duplicated definitions. For example, the following two definitions are 92% matching while the difference is only in the words in italics (“virtual representation” and “digital replica”):
@@
Table 2 shows a list of some definitions that were extracted from the dataset. The rest of definitions can be found on the GitHub repository [^6] https://github.com/xxxx/xxxx [will be added after publication].
Table 2: Ten examples of extracted DT definitions.
@@ 3.2 Digital Twin components identification
Data representation
Modeling Setup for modelling and analysis towards specific services, application use case and scenarios. Simulation models
@@
System aspects System and design characteristics of the digital twin, relating to integration, governance, security and privacy, and ethics. Data validation
@@
    DTs are envisioned to deal with multi-temporal and -scale heterogeneous data, which could be from multiple sources, either obtained from physical entities (e.g., as attribute data or dynamic condition data from sensors) or generated by virtual models (e.g., results from simulations within the DT). We generalize three related components under this category:
@@
      Real-time data: Real-time data updates (i.e., of the target entity as facilitated by sensors and Internet of Things) or dynamic data outputs from real-time simulation process.
@@
      Data representation: Formalized and standardized digital representation of data, including definitions for data models, schema, ontologies, and digitization methods.
@@
    System aspects
    Aspects of integration, governance, security and privacy, and data ethics are requirements for a DT system to coexist with other systems in the connected DT paradigm) (International Organization for Standardization, 2023), for example, matters related to data collection and access. Such system aspects could include (but are not limited to) the following:
@@ 4 Results
4.1 Component Analysis
We conducted component frequency and Chi-square analysis to examine the association between DT components (Table 3) and various domains, including urban, building, manufacturing, and architecture. The use of Chi-square test is particularly useful in our study due its ability to handle categorical data. The significance was tested against the null hypothesis $H_{0}$ (No significant contribution) and the alternative hypothesis $H_{1}$ (There is a significant contribution) for each component. The chi-square statistic of 109.76 and a p-value of 2.48e-07indicated a significant relationship between the components and their respective domains. We calculated the Standardized residuals to identify specific deviations, with significant residuals (absolute values greater than 2) highlighting components that were either over-represented or underrepresented. Notably, “IoT sensors” and “Policy” were significantly more frequent in the urban domain, while “2D/3D data” showed higher frequencies in the building and architecture domains but lower in manufacturing. This analysis, including adjustments for zero and low values, provided a detailed understanding of the distribution and significance of digital twin components across different fields, offering critical insights for the development and application of DT in various sectors (Figures 8 and 9). We also performed a correlation and dendrogram clustering analysis between all component residuals to explore which components are joined in terms of significance and insignificance across all domains (Figure 11) and for the built environment (Figure 12). The analysis clearly divides these components into two big clusters of correlations.
Cloud computing, HPC, Security protocols, Real time data, simulation models, and AI/ML models are highly correlated with each other and can be segregated into one group of components. On the other hand, Data representation, Data validation, Visualization, and Policy are highly correlated with each other and can be segregated into another group of components. We refer to the first group of components as HPRT (High-Performance Real-Time) Digital Twins. which is characterized by real-time big-data dynamic and adaptable components, while the second group can be referred to as LTDS (Long Term Decision Support Digital Twins), which are characterized by near-to-far real-time data management and visualization components. The first set of components is used in computationally intensive applications, such as manufacturing, self-driving cars, and aerospace, while the second set of components is used in applications that require long-term decision support, such as urban planning and building management.
@@ 4.1 Component Analysis
In the following sub-sections, we provide comprehensive analytical insight focusing on each domain.
Figure 8: Component frequency distribution across different domains.
Figure 9: Chi-square frequency distribution heatmap across different domains.
@@
Figure 12: Clustered correlation between components residuals across the built environment domains (not including manufacturing).
4.1.1 Building domain
@@
Figure 13: Components for (a) Building domain, (b) Architecture domain, and (c) Urban domain sorted by significance
@@
AI/ML models (residual of -1.5) and Real-time data (residual of -1.4) are less significant in the architecture domain. This might be due to the lack of data and the complexity of the data in the architecture domain. The integration of BIM and real-time data streams from IoT sensors is complex, interoperability between different data schemas, sources, software and platforms is difficult, in addition to scalability issues (Lu et al., 2020; Lei et al., 2023a).
@@
4.1.2 Urban domain
Urban terms in digital twins (UDTs) reveal various insights and challenges. Our analysis shows that there are fewer significant components in UDT given the complexity of the domain and the requirements of massive computational and storage resources, among other challenges that hindered the adoption of some components, such as real-time data and AI/ML in the urban domain. In contrast, IoT sensors, data representation, data validation, and plicy-driven decision-making show greater significance (Figure 13 - c).
@@ 4.1.2 Urban domain
4.2 Definition Analysis
In this section, we compare the definitions from the existing dataset (800 definitions) and the expert survey to the components analysis for each domain. We used Large Language Models (LLMs) to extract the components from these definitions and compare them with the components identified through frequency analysis. Our analysis reveals some divergences from the generalized definitions of digital twins. For example, when focusing on the urban/city domains, real-time data, and AI/ML models are not a significant portion of the definitions (Figure 14). Also, given our analysis results, we identified minimum level of maturity requirements for each domain based on the CITYSTEPS framework by Haraguchi et al. (2024).
The following subsections present our findings from the LLM analysis of the dataset, highlighting the top qualifying definitions that appeared in the literature for each domain and our derived definitions.
@@ 4.2.1 Building and Architecture Digital Twin (BDT and ADT) definitions
The following is a breakdown to the definition words according to the Cambridge Dictionary:
  • 1.
    virtual: ‘created by computer technology and appearing to exist but not exist in the physical world’.
  • 2.
    representation: ‘the way that someone or something is shown or described’.
  • 3.
    physical: ‘relating to things you can see or touch or relating to the laws of nature’.
  • 4.
    asset: ‘something having value, such as a possession or property, that is owned by a person, business, or organization’.
  • 5.
    system: ‘a set of connected things or devices that operate together’.
  • 6.
    collaborative: ‘involving two or more people or organizations working together for a particular purpose’.
  • 7.
    static information: ‘data that does not change frequently or remains constant over time’.
  • 8.
    dynamic information: ‘information that is periodically updated, meaning it changes asynchronously over time as new information becomes available’.
  • 9.
    adapt: ‘to adjust to different conditions or uses or to change to meet different situations’.
Gathering insights from our experts, we highlighted “accurate 3D geometry of buildings” and “rich semantic information” as critical components in the BE DTs. Such findings align with our extract definitions from the developed corpus of research articles. According to the CITYSTEPS framework, for a BDT/ADT to be considered a true digital twin, it must reach at least maturity level 4 i.e., the 3D dynamic stage (Figure 15). This dynamic aspect can range from real-time updates to long-term changes (Figure 14). Therefore, a BIM model, whether hosted online or on local servers, that lacks regular updates reflecting changes in the real-world entity does not meet the criteria for a digital twin.
@@ 6 Conclusion
The findings revealed significant variations in digital twin descriptions and paradigms across different domains, highlighting the complexity and evolving nature of this concept. We identified two primary categories of digital twins, and we found that all the built-environment-related DTs fit on the first one:
  1. 1.
    Long Term Decision Support (LTDS) Digital Twins: These digital twins focus on data representation, validation, and policy-driven decision-making, emphasizing long-term planning and management over real-time data integration. They are prevalent in building, architecture, and urban digital twins.
  2. 2.
    High-Performance Real Time (HPRT) Digital Twins: These digital twins are associated with real-time data processing, AI/ML models, and simulation capabilities, primarily used in computationally intensive applications outside the built environment.
Based on the extracted definitions and component analysis, and we deduced the following definitions for Building/Architecture Digital Twins (BDT/ADT) and Urban/City Digital Twins (UDT/CDT):
  1. 1.
  2. 2.
    A City/Urban Digital Twin is a spatial-temporal virtual representation of a real-world urban area or city, mirroring its states during its lifecycle through IoT sensors. It is used to monitor and analyze urban systems across different time spans to aid in decision-making and can be extended to simulate and predict various states and scenarios. It can consists of multiple BDTs or ADTs.
Although this research addresses the current definition of Digital Twins (DTs), future trends indicate that some components are becoming increasingly prominent such as AI/ML, HPC, and Real-time data adoption. Thus, future research is needed to redefine DTs according to new trends, techniques, and tools. This reflects the fact that Building, Urban, and City DTs have not yet reached their state of equilibrium regarding a fully indicative definition due to their complexity and ever-changing landscapes. This work marks the first step towards a consensus.

### v2

@@ 3.1 Sample Development
This study is built on two main data sources. The first consists of 1) full-text articles, while the second involves a 2) expert survey which is used to validate the results from a Built Environment expert perspective. From these two datasets, we have derived: the 3) DT definitions dataset and the 4) DT components dataset. In this section, we overview our methodology of collecting the sample data and extracting the definitions’ and components’ datasets that are used for analysis. We keep our methodology as detailed as possible for the purpose of reproducibility and reusability across different domains and studies.
@@ 3.1.1 Full-text articles dataset
Full-text samples were acquired using an Application Programming Interface (API) provided by Elsevier [^5] https://dev.elsevier.com/, which provides interfaces to Scopus, Science Direct, and other Elsevier products. We chose to focus on the Elsevier API because it is one of the largest publishers of scientific literature and provides a comprehensive database of peer-reviewed literature across myriads of fields, thus well capturing the body of knowledge without bias. We used the API to search for articles containing the following terms: “Digital Twin”, “City Digital Twin”, “Urban Digital Twin”, “Smart City”, and “Building Digital Twin” in the title, abstract, or keywords. The search was conducted in English and covered articles published between 2000 and 2024. The search returned a total of 15,353(Figure 6). The dataset was also collected from various sources (Figure 5), notably the top are from manufacturing (e.g, IFAC Journal of Systems and Control, Journal of Manufacturing Systems, Procedia Manufacturing); and the Built Environment (e.g, Automation in Construction, Transportation Research Procedia, Journal of Building Engineering, Energy and Buildings, and Building and Environment).
@@
Figure 6: Distribution of subjects in the selected dataset. The total search returned 15,353 results; however, this figure represents a subset of articles with valid subject metadata relevant to the analysis after cleaning the data.
@@ 3.1.2 The expert survey dataset
The expert survey invited a structured panel of experts, who have domain expertise in DT related to the Built Environment, aiming to generate a consensus on the definition of DTs. A total of 52 international experts from different sectors, such as government, industry (e.g. companies, consultancies), and non-government organizations (NGOs), were selected to build the panel, contributing with diverse backgrounds and experiences in this topic. In particular, the panel of domain experts represents 23 countries, with the majority working in Europe. Of these participants, 14 were from industry, 12 were from government, and the remaining 23 were affiliated with universities or research institutes. Notably, 44% of respondents reported over 20 years of experience in fields integral to digital twins. Through an online questionnaire, all participants were asked to identify a digital twin based on their understanding and practice.
@@ 3.1.3 The definitions dataset
This pattern extracts all sentences that match “digital twin/s (is/are/can be/could be) (defined as/described as/ characterized by) …”. Many of the extracted definitions are valid, and some are not valid definitions such as: “Digital Twin is a revolutionizing technology in the industry 4.0 era” and “Digital Twin is defined as a new paradigm in simulation”. Therefore, we applied NLP and LLMs to filter the dataset, aiming to address two main issues: Incomplete and Fuzzy Duplicated definitions.
@@
To perform the filter, we first excluded all the incomplete definitions using LLMs, namely, the Generative Pre-trained Transformer (GPT4) model. Then, we manually validated a subset of the results.
Secondly, to overcome the FDD issue, we first clustered the definitions using sentence embeddings Reimers and Gurevych (2019) and k-Means clustering into three iterations of clustering 400, 100, and 50 clusters. The Sentence Embeddings first convert all the definitions into a lower dimensional vector where the position of each definition in the vector space represents its semantic meaning. By doing this, the closer the definitions in the vector space, the more semantically similar these sentences are to each other. Each sentence is represented as a feature vector $\vec{X}_{d}=[x_{0},x_{1},x_{2},\ldots,x_{n}]$ where $X_{d}$ is the definition sentence vector. We used the cosine similarity distance to measure the similarity between different definitions in the definition dataset. The cosine similarity distance between two vectors $\mathbf{X}$ and $\mathbf{Y}$ is given by:
@@
After clustering the definitions into groups, we applied the Fuzzy Matching algorithm within each group to figure out how each pair of definitions matches. By doing so, we excluded semi-duplicated definitions. For example, the following two definitions are 92% matching while the difference is only in the words in italics (“virtual representation” and “digital replica”):
@@
Table 2 shows a list of some definitions that were extracted from the dataset. The rest of definitions can be found on the GitHub repository [^6] https://github.com/ualsg/dt-definitions.
Table 2: Ten examples of extracted DT definitions. The rest of definitions can be found on the GitHub repository
@@ 3.2 Digital Twin components identification
Data modeling
Analysis and services Setup for modelling and analysis towards specific services, application use case and scenarios. Simulation models
@@
System governance System and design characteristics of the digital twin, relating to integration, governance, security and privacy, and ethics. Data validation
@@
    DTs are envisioned to deal with multi-temporal and -scale heterogeneous data, which could be from multiple sources, either obtained from physical entities (e.g., as attribute data or dynamic condition data from sensors) or generated by virtual models (e.g., results from simulations within the DT) in both directions (Bi-directional) either from the physical twin to the virtual twin or vice versa using actuators or using human in the loop. We generalize three related components under this category:
@@
      Real-time data: Real-time data updates (i.e., of the target entity as facilitated by sensors and Internet of Things) or dynamic data outputs from real-time simulation process. Also, this includes data that are transferred the other direction, i.e, from the virtual twin to the physical twin via actuators or human-informed-decision in the loop.
@@
      Data modeling: Formalized and standardized digital representation of data, including definitions for data models, schema, ontologies, and digitization methods.
@@
    System governance
    Aspects of integration, governance, security and privacy, and data ethics are requirements for a DT system to coexist with other DT systems) (International Organization for Standardization, 2023), for example, matters related to data collection and access. Such system aspects could include (but are not limited to) the following:
@@
3.3 Component Identification and Analysis
We conducted component frequency and Chi-square analysis to examine the association between DT components (Table 3) and various domains, including urban, building, manufacturing, and architecture. The use of Chi-square test is particularly useful in our study due its ability to handle categorical data. The significance was tested against the null hypothesis $H_{0}$ (No significant contribution) and the alternative hypothesis $H_{1}$ (There is a significant contribution) for each component. The chi-square statistic of 109.76 and a p-value of 2.48e-07 indicated a significant relationship between the components and their respective domains. We calculated the Standardized residuals to identify specific deviations, with significant residuals (absolute values greater than 2) highlighting components that were either over-represented or underrepresented. Notably, “IoT sensors” and “Policy” were significantly more frequent in the urban domain, while “2D/3D data” showed higher frequencies in the building and architecture domains but lower in manufacturing. This analysis, including adjustments for zero and low values, provided a detailed understanding of the distribution and significance of digital twin components across different fields, offering critical insights for the development and application of DT in various sectors (Figures 8). We also performed a correlation and dendrogram clustering analysis between all component residuals to explore which components are joined in terms of significance and insignificance across all domains (Figure 10). The analysis clearly divides these components into two big clusters of correlations.
To extract the definitions across domains, we compared the definitions from the existing dataset (800 definitions) and the expert survey to the components analysis for each domain. LLM was used to extract the components from these definitions and compare them with the components identified through frequency analysis.
@@ 4 Results
4.1 Component Analysis Across Domains
Based on the Chi-square test and the correlation clustering of components, it is noticed that Cloud computing, HPC, Security protocols, Real time data, simulation models, and AI/ML models are highly correlated with each other and can be segregated into one group of components. On the other hand, Data representation, Data validation, Visualization, and Policy are highly correlated with each other and can be segregated into another group of components. We refer to the first group of components as HPRT (High-Performance Real-Time) Digital Twins, characterized by real-time, big-data-driven, dynamic, and adaptable components. The second group is referred to as LTDS (Long-Term Decision Support) Digital Twins, characterized by near-to-far real time data management and visualization components. The first group of components is used in computationally intensive applications, such as manufacturing, self-driving cars, and aerospace and rarely used in the BE and AEC domains. In contrast, the second group is suited for applications requiring long-term decision support, such as urban planning and building management.
@@ 4.1 Component Analysis Across Domains
In the following sub-sections, we provide comprehensive analytical insight and key results focusing on each domain  [^7] Although Building, Architecture, and Urban terms can be identified as different scales of the BE or the AEC rather than domains perse, they are segregated into different domains in this study because it is noted that there are some differences in the requirements and understanding of DT of each scale..
Figure 8: Chi-square frequency distribution heatmap across different domains and component frequency distribution across different domains.
@@
Building domain
@@
Figure 11: Components for the BE domains (a) Building, (b) Architecture, and (c) Urban sorted by significance. The manufacturing domain is excluded since its results is out of the scope of the research
@@
AI/ML models (residual of -1.5) and Real-time data (residual of -1.4) are less significant in the architecture domain. This might be due to the lack of data and the complexity of the data in the architecture domain.
@@
Urban domain
Urban terms in digital twins (UDTs) reveal various insights and challenges. Our analysis shows that there are fewer significant components in UDT given the complexity of the domain and the requirements of massive computational and storage resources, among other challenges that hindered the adoption of some components, such as real-time data and AI/ML in the urban domain. In contrast, IoT sensors, data representation, data validation, and policy-driven decision-making show greater significance (Figure 11 - c).
@@
4.2 Derived Definitions and Implications
In this section, we show the results of the derived definitions based on the frequency analysis. The results reveal some divergences from the generalized definitions of digital twins. For example, when focusing on the urban/city domains, real-time data, and AI/ML models are not a significant portion of the definitions (Figure 12). Also, given our analysis results, we identified minimum level of maturity requirements for each domain based on the CITYSTEPS framework by Haraguchi et al. (2024).
The following subsections present the findings from the LLM analysis of the dataset, highlighting the top qualifying definitions that appeared in the literature for each domain and our derived definitions.
@@ 4.2.1 Building and Architecture Digital Twin (BDT and ADT) definitions
Gathering insights from the experts’ responses, we highlighted “accurate 3D geometry of buildings” and “rich semantic information” as critical components in the BE DTs. Such findings align with the extract definitions from the developed corpus of research articles. According to the CITYSTEPS framework, for a BDT/ADT to be considered a true digital twin, it must reach at least maturity level 4 i.e., the 3D dynamic stage (Figure 13). This dynamic aspect can range from real-time updates to long-term changes (Figure 12). Therefore, a BIM model, whether hosted online or on local servers, that lacks regular updates reflecting changes in the real-world entity does not meet the criteria for a digital twin.
@@ 5 Discussion and Future Directions
Comparison with traditional approaches
Traditional approaches to analyzing definitions and concepts, such as manual literature reviews or expert-driven content analysis, often rely on smaller datasets due to the time-intensive nature of these methods. While these methods can provide nuanced insights, they are inherently limited in scalability and subject to potential bias introduced by individual reviewers.
Our approach leverages advanced NLP techniques, enabling us to process and analyze a large corpus of 15,000 articles efficiently. This methodology provides distinct advantages over traditional methods.
A key strength of this approach lies in its ability to derive definitions from a collective understanding of the concept. By analyzing a large and diverse set of papers discussing DTs across different disciplines, our methodology captures the broader agreement within the scientific community while identifying nuances and variations. This collective perspective enables the development of a definition that reflects a consensus-driven understanding, ensuring relevance and applicability across domains.
@@ 6 Conclusion
The findings revealed significant variations in digital twin descriptions and paradigms across different domains, highlighting the complexity and evolving nature of this concept. We identified two primary categories of digital twins, Long Term Decision Support (LTDS) DTs and High-Performance Real Time (HPRT) DTs. We found that all the built-environment-related DTs fit on the first one:
Although this research addresses the current definition of Digital Twins (DTs), future trends indicate that some components are becoming increasingly prominent such as AI/ML, HPC, and Real-time data adoption. Thus, future research is needed to redefine DTs according to new trends, techniques, and tools. This reflects the fact that Building, Urban, and City DTs have not yet reached their state of equilibrium regarding a fully indicative definition due to their complexity and ever-changing landscapes. This research offers a foundational framework for defining Digital Twins in the built environment and serves as a first step toward building consensus. The methodology and definitions presented here provide a sustainable approach that encourages researchers and practitioners to adopt these definitions in their implementations or refine them based on specific requirements. We encourage researchers to deploy the same methodology and revisit the definition periodically to track its evolution over time, reflecting advancements in technology and evolving understandings of the DT entity.
@@
CRediT author statement
MA: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Visualization; EM: Methodology, Validation, Formal analysis, Investigation, Writing - Original Draft; BL: Validation, Formal analysis, Investigation, Data Curation, Writing - Review & Editing; MQ: Software, Formal analysis, Writing - Review & Editing, Visualization; CM: Conceptualization, Writing - Review & Editing; FB: Conceptualization, Methodology, Investigation, Writing - Review & Editing, Visualization, Supervision, Project administration, Funding acquisition;